(4y/9) - (2y/3) = 10
Answer by Alan3354(69443)

(4y/9) - (2y/3) = 10
--------------
No question, no request.
If you want to solve for y, you should say, "Solve for y."
------------
(4y/9) - (6y/9) = 10
4y - 6y = 90
-2y = 90
y = -45